Translate the problems into a pair of linear equations in two variables. Solve equations using either elimination or substitution. State your answer for the specified variable.
A plane flies 400 miles with the wind and 300 miles against the wind in the same length of time. If the speed of the wind is 22 mph, what is the speed of the plane in still air?

Answer by checkley77(12844)

D=RT
400=(R+22)T WITH THE WIND.
400=RT+22T
300=(R-22)T AGAINST THE WIND.
300=RT-22T
400=RT+22T SUBTRACT
-----------------
-100=-44T
-44T=-100
T=-100/-44
T=2.27 HOURS FOR THE TRIP.
400=(R+22)2.27
400=2.27R+49.94
2.27R=400-49.94
2.27R=350.06
R=350.06/2.27
R=154.21 MPH FOR THE PLANE IN STILL AIR.
PROOF:
300=(154.21-22)*2.27
300=132.21*2.27
300=300